OpenStudy (anonymous):

In this expression below shown below, what number does the letter n represent? 875 divided by 5=(n+70+5)divided by 5

jimthompson5910 (jim_thompson5910):

875/5=(n+70+5)/5 875/5=(n+75)/5 875*5=5(n+75) 4375=5(n+75) 4375=5n+375 4375-375=5n 4000=5n 4000/5 = n 800 = n n = 800

OpenStudy (anonymous):

uh.............

jimthompson5910 (jim_thompson5910):

Another way to look at it 875 = 800 + 70 + 5 So if you erase the 800, and replace it with 'n', then we get 875 = n + 70 + 5 So n = 800 This is working backwards, which is often helpful when solving problems.

OpenStudy (anonymous):

I thought u \[875\div5 and \it equals 175 that =N\]

jimthompson5910 (jim_thompson5910):

You could do that, but you'll multiply that result by 5 again (to get rid of the 5 on the right side) So you'll be left with 875 = n + 70 + 5

OpenStudy (anonymous):

so, after u do the whole problem the result u get is wat n equals

jimthompson5910 (jim_thompson5910):

Exactly, after doing all that, the result is n = 800

OpenStudy (anonymous):

oooo i thought it was just do 875 divided by 5 and thats it

OpenStudy (anonymous):

Well, Thanks!! ^-^

jimthompson5910 (jim_thompson5910):

Well that's part of it, but I prefer to just deal with the numerators (since the denominators go away anyway). You're welcome